When price is Rs. 20 per unit, demand for a commodity is 500 units. As the price falls to Rs. 15 per unit, demand expands to 800 units.

Calculate elasticity of demand.

2. Mention type of elasticity

3. Type of relationship between price and quantity

1
Expert's answer
2020-07-13T14:04:19-0400

Change in Quantity ("\Delta"Q)=800-500=300 units

Change in price("\Delta"P)=15-20= -Rs.5

Elasticity Demand="\frac{ΔQ}{ΔP}× \frac{P}{Q}=\frac{300}{-5}×\frac{20}{500}= -2.4"


Demand is highly elastic as ED is greater than 1 i.e ED>1


There is inverse relationship between the price and quantity demanded because of the negative sign.